The derivation of f(tan x) w.r.t g(sec x) at x= pi/4 where f`(1)=2 and g`(sqrt2)=4 is

Grade:12th pass

1 Answers

Vinod Kumar
9 years ago
[d(f(tanx))/dx] / [d(g(secx))/dx]=[(f '(tan(x))*sec2x] / [(g '(secx)) secx*tanx]

put x=pi/4

=[(f '(1)*(v2)^2] / [(g'(v2))*v2*1]

=(2*2) / (4*v2)

=1/v2


Thanks and Regards,
Vinod Kumar,
askIITians faculty,
IIT- Delhi